Problem 2


The Senior High School researchers claim that 20% of Senior High School male students have tried smoking cigarette. After collecting 150 random samples, they found 40 of them have tried smoking cigarette.


Hypothesized proportion (p̂), th value of q, and finally the value of z.


*


p̂ = 0.27, q = 0.9, z = 2.05


p̂ = 0.27, q = 0.8, z = 2.14


p̂ = 0.27, q = 0.8, z = 2.0412


p̂ = 0.33, q = 0.7, z = 0.8

1
Expert's answer
2022-06-16T06:30:07-0400
"\hat{p}=\dfrac{x}{n}=\dfrac{40}{150}=\dfrac{4}{15}\approx0.27"

"q=1-p=1-0.2=0.8"

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{40/150-0.2}{\sqrt{\dfrac{0.2(1-0.2)}{150}}}=2.0412"



"p\u0302 = 0.27, q = 0.8, z = 2.0412"